Manuel invested his savings in two investment funds. The $12,000 that he invested in Fund A returned a 1% profit. The amount that he invested in Fund B returned a 5% profit. How much did he invest in Fund B, if both funds together returned a 2% profit?
Found 2 solutions by TimothyLamb, ewatrrr:
Answer by TimothyLamb(4379)

x = amount invested in A
y = amount invested in B
---
x + y = 12000
0.01x + 0.05y = 0.02*12000
---
put the system of linear equations into standard form
---
x + y = 12000
0.01x + 0.05y = 240
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = amount invested in A = 9000
y = amount invested in B = 3000
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by ewatrrr(24785)

x the Fund B amount
.05x + .01(12000 - x) = .02(12000)
x = .01(12000)/.04 = $3000